OpenStudy (anonymous):

need help fast! What is the slope of the line that passes through the points (1, 3) and (5, –2)?

OpenStudy (anonymous):

You need me fast?

OpenStudy (anonymous):

Just put your points in and solve

pooja195 (pooja195):

not y= its slope = :P

OpenStudy (anonymous):

Use the slope formula \( \huge m = \frac{Y_2- Y_1}{X_x-X_1} \)

OpenStudy (anonymous):

Yeah I seen that lol @pooja195

OpenStudy (anonymous):

I did and I cant get the answer and I need the answer

OpenStudy (anonymous):

Use the slope formula \( \huge m = \frac{3- (-2)}{1-5} \) m = slope just solve now

OpenStudy (anonymous):

Can you solve that above??

OpenStudy (anonymous):

3 - (-2) = what??

pooja195 (pooja195):

@Denali19 do yr own work

OpenStudy (anonymous):

Ok, I am off to get chicken nuggets from McD's :-)